A car accelerates from initial velocity of 10 m/s. The acceleration id increasing uniformly from zero to 8 m/s^2 in 6 secs. During the next two seconds, the car decelerates at a constant rate of 2 m/s^2. 1) Which of the following most nearly gives the velocity after 8 seconds. 2.) Which of the following most nearly gives the distance travelled after 6 seconds.


1
Expert's answer
2021-12-20T10:27:35-0500

1.

"v=v_0+\frac{\dot a_1 t_1^2}2-a_2t_2=30~\frac ms."

2.

"s=v_0t_1+\frac{\dot a_1t_1^3}6=108~\frac ms ."